OpenStudy (anonymous):

A rocket is fired upward from some initial distance above the ground. It's height in feet, h, above the ground, t seconds after it is fired is give by h=-16t^2+48t+4864. What is the rockets maximum height? How long does it take the rocket to reach it's maximum height? After it is fired, the rocket reaches the ground at t= __sec.

OpenStudy (anonymous):

do you know calculus?

OpenStudy (anonymous):

derivative to be specific

OpenStudy (anonymous):

not at all, just taking a test at the moment.

OpenStudy (anonymous):

Oh ok, then do you know how to find vertex of a parabola?

OpenStudy (anonymous):

i honestly know nothing...haha

OpenStudy (anonymous):

y = ax^2 + bx + c equation for vertex = \[\frac{ -b }{ 2a }\] In this case, a = -16 and b = 48 Plug it in the equation to find vertex The vertex of a downward parabola gives you the maximum height of the rocket.

OpenStudy (anonymous):

vertex = time of maximum height. 48/-(-16) = 3 seconds

OpenStudy (anonymous):

maximum height = -16(3)^2 + 48(3) + 4864

OpenStudy (anonymous):

so maximum height is 4864? & after it is fired how many secs till it reaches the ground?

OpenStudy (anonymous):

then to find when it reaches the groung again, set h = 0 then solve for the equation 0= -16t^2 + 48t + 4864 0= -16(t^2 - 3t - 304) 0 = -16(t-19)(t+16) t= 19, -16, since t cannot be negative, the only answer is 19 seconds for time on the ground

OpenStudy (anonymous):

I gave you the equation for maximum height. it's not 4864

OpenStudy (anonymous):

my time already ran up, but thank you away! :)